OpenStudy (ladylauren):

Will fan and medal A candle burns down at the rate of 0.5 inches per hour. The original height of the candle was 8 inches. Part A: Write a list of 6 ordered pairs to show the height of the candle in inches (y) as a function of time in hours (x) from the first hour after it started burning. For example, the point (0, 8) would represent a height of 8 inches after 0 hours. Explain how you obtained the ordered pairs.

OpenStudy (photon336):

You could represent this using a linear function y = m(x)+b where the height is initially 8 inches. if the candle is decreasing at a rate of 0.5 inches per hour. how do you think we could represent this?

OpenStudy (photon336):

@ladylauren

OpenStudy (ladylauren):

so you subtract 0.5 every hour

OpenStudy (photon336):

well look at it this way. each hour we know that the height is decreasing by 0.5 inches. so after x amount of hours the height decreases by 0.5(x) follow?

OpenStudy (ladylauren):

yes

OpenStudy (photon336):

but we also know that the height can't keep decreasing forever. it's going to stop after a while once it reaches 8 inches. but we know that the height is initially 8 inches. y = 8-0.5(x)

OpenStudy (ladylauren):

so we subtract 0.5 6 times?

OpenStudy (photon336):

On the y axis is the height of the candle, on the x axis it's the amount of hours.

OpenStudy (photon336):

what i'm saying here is that we can build a function that represents the candle height and time

OpenStudy (ladylauren):

ok

OpenStudy (photon336):

Now for your ordered pairs you can plug in any x value into this equation starting from x = 0 and then plot the points. \[y_{1} = 8-0.5x \]

OpenStudy (ladylauren):

ok so it would be 7.5

OpenStudy (ladylauren):

after that

OpenStudy (photon336):

have you learned about linear equations like in the form of a line?

OpenStudy (ladylauren):

yes

OpenStudy (photon336):

now you have a function that represents the height of the candle as a function of time. now let's find the height of the candle at 0 hours so in zero hours right, we can plug in zero into our equation. \[8-(0)*\frac{ 1 }{ 2 } = 8~inches \] we get 8 inches. this makes sense. now what about after an hour? we plug in 1 where we see x and get 7.5 inches. \[8-(1)*\frac{ 1 }{ 2 } = 7.5~inches \] Here are our points so far (0,8) (1,7.5) can you see the pattern and continue from there?

OpenStudy (ladylauren):

so it is (0.8),(1,7.5),(2,7),(3,6.5),(4,6),(5,5.5)

OpenStudy (photon336):

Yes, and all the information about this is contained in our equation

OpenStudy (ladylauren):

ok

OpenStudy (photon336):

what do you notice about the height after each hour just by looking at the points.

OpenStudy (ladylauren):

can you help me with part 2

OpenStudy (ladylauren):

it decreases 0.5 every hour and they are on a straight line so it is liner

OpenStudy (photon336):

yes. I'll give you a hint but not the answer

OpenStudy (ladylauren):

ok

OpenStudy (ladylauren):

Is this relation a function? Justify your answer using the list of ordered pairs you created in Part A.

OpenStudy (photon336):

What do you think?

OpenStudy (ladylauren):

i think it would be a linear function

OpenStudy (photon336):

okay, but why?

OpenStudy (ladylauren):

because when i put it on a graph it is in a straight line

OpenStudy (photon336):

that's one reason

OpenStudy (photon336):

the other reason is that in every function three must be one only x value for every y value

OpenStudy (ladylauren):

ok like in function 1 input for each output

OpenStudy (photon336):

another way we can actually test this by placing a vertical line through our function like this. can you see why one is a function and not the other? 

OpenStudy (ladylauren):

cause the line on the function is straight and the other one isent

OpenStudy (photon336):

Well, yes kind of. if you notice that for the circle the line passes through two points. one x value has two y values It can't be a function .

OpenStudy (ladylauren):

oh i see now

OpenStudy (ladylauren):

thank you so much this was realy helpfull

OpenStudy (photon336):

no problem